OpenStudy (anonymous):

The lesson is about completing the square. help. Solve: x^2 - 4/3x = 4/9 = 0

OpenStudy (anonymous):

Do you have to answer is fractions?

OpenStudy (mathstudent55):

You have two equal signs.

OpenStudy (anonymous):

should read - 4/9 not = 4/9

OpenStudy (mathstudent55):

\(x^2 - \dfrac{4}{3}x - \dfrac{4}{9} = 0\)

OpenStudy (mathstudent55):

First step, move constant term to right side. Add 4/9 to both sides.

OpenStudy (mathstudent55):

\(x^2 - \dfrac{4}{3}x - \dfrac{4}{9} = 0\) \(x^2 - \dfrac{4}{3}x ~~~~~~~~~~~= \dfrac{4}{9}\)

OpenStudy (mathstudent55):

Since the coefficient of the x^2 term is one, we can do the next step.

OpenStudy (mathstudent55):

Take the coefficient of the x-term, -4/3. Divide it by 2. Then square it. Then add it to both sides.

OpenStudy (anonymous):

Bring everything over to one side assuming the 1st equals is a plus..... x^2 - 4/3x + 4/9 = 0 Add and subtract the quantity (b/2)^2 or ((4/3)/2)^2 x^2 - 4/3x - 4/9 + 4/9 - 4/9 = 0 Then you can make a perfect square with the "x^2 - 4/3x + 4/9" -4/9 - 4/9 = 0 Making the perfect square you take the "b" term and divide it by two and make that the constant term inside the quantity squared. (x-2/3)^2 - 8/9 = 0

OpenStudy (mathstudent55):

\(x^2 - \dfrac{4}{3}x ~~~~~~~~~~~= \dfrac{4}{9}\) Half of -4/3 is -4/6 = -2/3. Squaring -2/3 = 4/9 Now we add 4/9 to both sides. \(x^2 - \dfrac{4}{3}x + \dfrac{4}{9} = \dfrac{4}{9} + \dfrac{4}{9}\) The left side is now the square of a binomial: \( \left( x - \dfrac{2}{3} \right)^2 = \dfrac{8}{9} \) Now you can solve the equation by taking the square root of both sides. Remember that if \(x^2 = a\), then \( x = \pm \sqrt{a} \)